OpenStudy (anonymous):

48xy^4 - 72xy^2 ------------- < divided line 12xy^2 A. 4xy2 - 6xy2 B. 4y2 - 6 C. 4xy - 6y D. 4y - 6

OpenStudy (anonymous):

\[\frac{48xy^4-72xy^2}{12xy^2}\]Treat the coefficients and variables separately. Let's do the coefficients first. You're just going to look at the 48, -72, and 12 first.

OpenStudy (anonymous):

What's the greatest common factor for those three numbers?

OpenStudy (anonymous):

72 is not negative ? but is 72 the greatest common factor ?

OpenStudy (anonymous):

A common factor is a number that will divide into all three numbers. What can you divide 48, 72, and 12 by?

OpenStudy (anonymous):

No lie i don't even know ?

OpenStudy (anonymous):

No worries! For example, 2 goes evenly into 48, 72, and 12. Same thing with 3.

OpenStudy (anonymous):

So the way to start these problems is to look at what you can "pull out" of all the numbers.

OpenStudy (anonymous):

so 3 &2 both go into 48,72,12 evenly ?

OpenStudy (anonymous):

Exactly. So does 6. If we divide each number by 6, we get 8 (from 48), 12 (from 72), and 2 (from 12). With me so far?

OpenStudy (anonymous):

so you got 8 outta 48 and 12 outta 72 & 2 outta 12 ?

OpenStudy (anonymous):

Right. So here's what this means. Instead of writing \[\frac{48xy^4-72xy^2}{12xy^2}\]you could write \[\frac{8xy^4-12xy^2}{2xy^2}\]which definitely looks a little nicer

OpenStudy (anonymous):

Now what do i do from there ?

OpenStudy (anonymous):

Now, the numbers aren't completely reduced. What's the common denominator for 8, 12, and 2?

OpenStudy (anonymous):

4 goes in to 8 & 6 goes into 12 and 2 goes in to 2 ?

OpenStudy (anonymous):

Those divide evenly into the numbers, but what you're looking for is a single number that goes into all of them. 2 goes into all three of the numbers, so that's the number you're looking for. Even though 4 goes into 8, 4 doesn't go into 2. When you pull out a 2 from each one, you get \[\frac{4xy^4-6xy^2}{1xy^2}\] Now let's look back at the original problem. It was possible to pull out a 12 at the very beginning, which would have gotten us here in one step instead of two steps. 12 was the largest number that would have gone into 48, 72, and 12. That means that 12 is our least common denominator. Let me know when you're ready to go on.

OpenStudy (anonymous):

Oops, I typed "least common denominator." What I meant was "greatest common factor."

OpenStudy (anonymous):

Okay im ready

OpenStudy (anonymous):

Alright. Now we're going to ignore the original problem. All we care about now is this: \[\frac{4xy^4-6xy^2}{xy^2}\]because we made it easier to deal with. Now we're going to mess with the variables.

OpenStudy (anonymous):

You're concerned now with the xy^4, xy^2, and xy^2. Treat the x's separately from the y's. Each term has one x. We can basically just cancel all of them out. Now we have \[\frac{4y^4-6y^2}{y^2}\] Okay with that?

OpenStudy (anonymous):

yeah so the answer is 4y2-6

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thank's alot

OpenStudy (anonymous):

No problem. Good luck.

OpenStudy (anonymous):

Thank's lol ill be pack tomorrow with like 90 more questions

OpenStudy (anonymous):

Back *

OpenStudy (anonymous):

Ew, sorry about that.